OpenStudy (anonymous):

What is the experimental probability that the next person you survey does NOT have a gaming system. You are writting an article on gaming systems for you school newspaper. You take a survey of 250 people ages 13 - 18 and ask whether they have a home gaming system, a portable gaming system, or no gaming system. Age 13 14 15 16 17 18 Home Gaming 12 22 26 32 24 26 Portable Gaming 5 5 4 9 7 6 No Gaming 18 12 8 6 12 16

jimthompson5910 (jim_thompson5910):

There are 250 people total Out of these 250 people there are 18+12+8+6+12+16 = 72 people who do not have a gaming system (added up the values in the bottom row) so the probability is 72/250 = 36/125 if you need it as a decimal number, then use a calculator to get 36/125 = 0.288, and this is a 28.8% chance

jimthompson5910 (jim_thompson5910):

btw for some very strange reason, all of the values in that table add to 240 (not 250) but the problem clearly states that "You take a survey of 250 people"

OpenStudy (anonymous):

O thanks!!! ya i realized that but its out of a book.. so! not really anything i can change there! lol

jimthompson5910 (jim_thompson5910):

yeah hard typo to catch unless you add up all those numbers (which many aren't willing to do)

OpenStudy (anonymous):

Ya! i ended up doing that and noticed it! but not gonna say anything! lol

jimthompson5910 (jim_thompson5910):

lol it might help to bring it up with your teacher and s/he might know what to do

OpenStudy (anonymous):

Haha thanks online school it is harder to do that ! but i am sure it will be fine :) Thanks for the help! btw!

OpenStudy (anonymous):

O you know this is my fault the top numbers for home gaming the last one should be 26 not 16 :)

jimthompson5910 (jim_thompson5910):

you're not able to contact the teacher?

jimthompson5910 (jim_thompson5910):

oh that accounts for the missing 10 that went AWOL, lol thanks

OpenStudy (anonymous):

Haha totally my fault i looked mack in the book and i typed it in wrong! My bad np!

jimthompson5910 (jim_thompson5910):

that's ok, there's a lot to keep track of and it happens all the time to me

OpenStudy (anonymous):

Haha ya forsure!! kk welll thanks allot i have like 2 finals to do now and again thanks for the help! :)

jimthompson5910 (jim_thompson5910):

good luck and yw

OpenStudy (anonymous):

Haha thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

The table numbers add to 250.

OpenStudy (anonymous):

@jim_thompson5910 you added incorrectly